Shift report 20011010 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Oct 10 2001 - 00:08:24 EDT

  • Next message: Hironori Ito: "new dndeta figures! (still BETA)"

    ------------------------------------------------------------
    Date: Wed Oct 10 00:08:24 2001, Shift: 16:00-24:00
    Supervisor: J.I.Jordre, with: 
    Log book#9 pages 131 - 133
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam stored and cogged at 21:20, still running. Started off with
       low ZDC rates, but towards the end of run 5274 MCR managed to 
       increase it by steering.
       (By the way, why does MCR always steer in BRAHMS' IR last. Couldn't
       they rather do STAR last since STAR doesn't want that high rates
       anyhow? Just a suggestion.) 
    2. Run 5273	aborted after 10mins when I found I had forgotten
     		to include FB3
       Run 5274	264.8k trigger 1 events
       Run 5276	~130k trigger 1 events @ midnight
    3. No new detector problems
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Oct 10 2001 - 00:08:50 EDT